OpenStudy (anonymous):

Find the minimum or maximum value of the function. Describe the domain and range of the function and where the function in decreasing or increasing: y=6x^2-1 I'm not sure how to figure out how to find the x and y values because its not already in standard form for me to find the vertex? What am I doing wrong?

OpenStudy (aum):

Vertex form is: y - k = a(x - h)^2 where (h,k) is the vertex. Rearrange y = 6x^2 - 1 to look like the equation above. y + 1 = 6(x - 0)^2 what is (h, k)?

OpenStudy (aum):

y + 1 = 6(x - 0)^2 y - (-1) = 6(x - 0)^2 y - k = a(x - h)^2 What is (h, k) ?

OpenStudy (aum):

If you are not familiar with the vertex form of a parabola you can skip the above method and calculate the x-coordinate of the vertex from the standard form of a parabola. For y = ax^2 + bx + c The x-coordinate of the vertex is -b/(2a) Here y = 6x^2 - 1 which can be written as follows: y = 6x^2 + 0x - 1 where a = 6, b = 0, c = -1 The x-coordinate of the vertex is -b/(2a) = -0/(2*6) = 0 put x = 0 and find y: y = 0 + 0 - 1 = -1 Vertex is at (0, -1)